OpenStudy (anonymous):

(5) to the third power x equals 5 to the the 12th what is x

OpenStudy (anonymous):

\[x^3=5^{12}\]\[x=5^{\frac{12}{3}}=5^4\]

OpenStudy (anonymous):

\[\sqrt[3]{x^3}=\sqrt[3]{5^{12}}=5^{12/3}=5^4\] We don't have to worry about the sign of the root because it's an odd numbered root so there's only one root not 2 (even power roots usually have a (+) and (-) solution)